OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 5| + 2 = 5 Step 2: |x – 5| = 5 – 2 Step 3: |x – 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = -3 -x – 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (amoodarya):

x-5= +-3 it means x-5=3 or x-5=-3 so 3rd

OpenStudy (anonymous):

thank you your the best

OpenStudy (amoodarya):

your welcome